# Dynamic Multimodal Prototype Learning in Vision-Language Models

### v1

@@ 4.1 Setup
CLIP [38] 59.81 23.24 52.91 60.72 35.48 46.43 43.09
CoOp [58] 63.33 23.06 55.40 56.60 34.67 46.61 42.43
CoCoOp [57] 62.81 23.32 55.72 57.74 34.48 46.81 42.82
CuPL [36] 61.45 23.86 54.61 61.02 35.13 47.21 43.65
VisDes [30] 59.68 21.17 52.36 57.16 33.78 44.83 41.11
TPT [41] 60.74 26.67 54.70 59.11 35.09 47.26 43.89
DiffTPT [11] 60.80 31.06 55.80 58.80 37.10 48.71 45.69
TDA [21] 61.35 30.29 55.54 62.58 38.12 49.58 46.63
DOTA [14] 61.82 30.81 55.27 62.81 37.52 49.64 46.60
DPE [55] 63.41 30.15 56.72 63.72 40.03 50.80 47.65
AWT [65] 63.01 30.93 56.31 63.06 39.71 50.60 47.50
RN50 ProtoMM 63.76 30.56 57.28 63.95 41.05 51.32 48.21
CLIP [38] 68.34 49.89 61.88 77.65 48.24 61.20 59.42
CoOp [58] 71.51 49.71 64.20 75.21 47.99 61.72 59.28
CoCoOp [57] 71.02 50.63 64.07 76.18 48.75 62.13 59.91
CuPL [36] 69.62 50.72 63.27 77.05 49.02 61.93 60.01
VisDes [30] 68.55 49.07 61.80 75.13 47.97 60.50 58.49
TPT [41] 68.98 54.77 63.45 77.06 47.94 62.44 60.81
DiffTPT [11] 70.30 55.68 65.10 75.00 46.80 62.28 60.52
TDA [21] 69.51 60.11 64.67 80.24 50.54 65.01 63.89
DOTA [14] 70.68 61.19 64.41 81.17 51.33 65.75 64.52
DPE [55] 71.91 59.63 65.44 80.44 52.26 65.93 64.44
AWT [65] 71.26 60.33 65.05 80.38 51.60 65.72 64.34
ViT-B/16 ProtoMM 72.01 64.02 65.93 80.87 51.97 66.96 65.69
[…]
@@ 4.2 Main Results
We compare our methods with state-of-the-art (SOTA) methods across 11 widely used classification datasets, using RN50 and ViT-B/16 backbones. The results presented in Table 1 highlight the superior performance of our method which consistently outperforms AWT [65] across all datasets under RN50 excluding Pets, with an averaged accuracy gain of 0.93%. When compared to TDA [21] and DOTA [14] using RN50, our method achieves the 2.04% and 0.94% averaged accuracy improvements, which is significant in test-time scenarios. In the Caltech101 [10] dataset, DPE [55] exhibits the highest performance, which is due to tuning the adapters for each testing sample. However, this process is more computationally and time-costing. For the ViT-B/16 backbone, we observe similar results, with our method surpassing the existing methods across most datasets, showing an average accuracy improvement of 1.08% over DPE. Especially on the fine-grained Flowers102 [33] datasaet, our method outperforms the second-best AWT by over 2.3%. We also evaluate our method with SOTA methods on the ImageNet [9] and its 4 out-of-distribution (OOD) variants, as shown in Table 2. The experimental results demonstrate that our method exhibits competitive generalization performance across all methods. Specifically, our ProtoMM outperforms TDA by 2.58% and AWT by 0.71% on average when using the RN50 backbone. Moreover, with the ViT-B/16 backbone, our method surpasses the second-best method, DOTA, by an average of 1.17% on the OOD dataset. Even compared to the training-based DPE, it achieves an average improvement of 1.03%. These results demonstrate the adaptability of our method to varied distributions. Notably, this generalization is achieved without any test-time optimization or tuning.
@@ 4.3 Ablation Study
Effectiveness of using multimodal prototypes. We examine the contributions of the distributed feature construction module and the multimodal prototype learning module, respectively. The results are summarized in Table 3. Row (1) presents the zero-shot CLIP [38] performance as a baseline. In Row (2), Eq. (5) stands for the distributed feature construction, which expands the textual prototypes by incorporating additional features with their importance weights, achieving 1% gains over CLIP on ImageNet with RN50. Row (3) shows the results of using both modules in our method, outperforming the variant that only uses Eq. (5) by over 1% on Caltech101 with ViT-B/16. These improvements indicate that dynamically incorporating multimodal prototypes leads to better representations and improved test-time adaptation.
@@
Comparison of Inference Time Our method is training-free, unlike methods that require prompt tuning or adapter tuning, such as TPT, DiffTPT, and DPE, which involve backpropagating through an expensive encoder or adapter during optimization. We evaluate the inference time using a single NVIDIA 3090 GPU. The experimental results in Table 4 indicate that our method is faster than the tuning-based methods. Although our method takes slightly more time than TDA, we achieve a notable performance improvement, outperforming TDA by over 2.4%. It is worth noting that the Sinkhorn algorithm [8] supports efficient parallel computation, allowing the distance between one distribution and multiple others to be evaluated simultaneously, as demonstrated in prior works [4, 65]. This parallelism further contributes to the comparable runtime of our method relative to TDA.
@@
CLIP [38] 13min 59.81 0
TPT [41] 14h 30min 60.74 +0.93
DiffTPT [11] 37h 25min 60.80 +0.99
TDA [21] 17min 61.35 +1.54
DOTA [14] 18min 61.82 +2.01
DPE [55] 2h 40min 63.41 +3.60
AWT [65] 21min 63.01 +3.20
Proto 21min 63.76 +3.95

### v2

@@ 4.1 Setup
CLIP [CLIP] 59.81 23.24 52.91 60.72 35.48 46.43 43.09
CoOp [coop] 63.33 23.06 55.40 56.60 34.67 46.61 42.43
CoCoOp [cocoop] 62.81 23.32 55.72 57.74 34.48 46.81 42.82
CuPL [CuPL] 61.45 23.86 54.61 61.02 35.13 47.21 43.65
VisDes [VisDesc] 59.68 21.17 52.36 57.16 33.78 44.83 41.11
TPT [TPT] 60.74 26.67 54.70 59.11 35.09 47.26 43.89
DiffTPT [DiffTPT] 60.80 31.06 55.80 58.80 37.10 48.71 45.69
TDA [TDA] 61.35 30.29 55.54 62.58 38.12 49.58 46.63
DOTA [dota] 61.82 30.81 55.27 62.81 37.52 49.64 46.60
DPE [DPE] 63.41 30.15 56.72 63.72 40.03 50.80 47.65
AWT [AWT] 63.01 30.93 56.31 63.06 39.71 50.60 47.50
RN50
ProtoMM 63.76 30.56 57.28 63.95 41.05 51.32 48.21
CLIP [CLIP] 68.34 49.89 61.88 77.65 48.24 61.20 59.42
CoOp [coop] 71.51 49.71 64.20 75.21 47.99 61.72 59.28
CoCoOp [cocoop] 71.02 50.63 64.07 76.18 48.75 62.13 59.91
CuPL [CuPL] 69.62 50.72 63.27 77.05 49.02 61.93 60.01
VisDes [VisDesc] 68.55 49.07 61.80 75.13 47.97 60.50 58.49
TPT [TPT] 68.98 54.77 63.45 77.06 47.94 62.44 60.81
DiffTPT [DiffTPT] 70.30 55.68 65.10 75.00 46.80 62.28 60.52
TDA [TDA] 69.51 60.11 64.67 80.24 50.54 65.01 63.89
DOTA [dota] 70.68 61.19 64.41 81.17 51.33 65.75 64.52
DPE [DPE] 71.91 59.63 65.44 80.44 52.26 65.93 64.44
AWT [AWT] 71.26 60.33 65.05 80.38 51.60 65.72 64.34
ViT-B/16
ProtoMM 72.01 64.02 65.93 80.87 51.97 66.96 65.69
Table 2: Performance comparison of accuracy ($\%$) on the out-of-distribution benchmark. All the compared methods are built upon CLIP-RN50 or CLIP-ViT-B/16 backbone. The two evaluation metrics Average and OOD Average are calculated by taking the mean accuracy across all five datasets and four OOD datasets excluding ImageNet. The best and second best results are denoted in bold and underline, respectively. refers to our implementation using their publicly released code.
@@ 4.2 Main Results
We compare our methods with state-of-the-art (SOTA) methods across 11 widely used classification datasets, using RN50 and ViT-B/16 backbones. The results presented in Table 1 highlight the superior performance of our method which consistently outperforms AWT [AWT] across all datasets under RN50 excluding Pets, with an averaged accuracy gain of 0.93%. When compared to TDA [TDA] and DOTA [dota] using RN50, our method achieves the 2.04% and 0.94% averaged accuracy improvements, which is significant in test-time scenarios. In the Caltech101 [Caltech] dataset, DPE [DPE] exhibits the highest performance, which is due to tuning the adapters for each testing sample. However, this process is more computationally and time-costing. For the ViT-B/16 backbone, we observe similar results, with our method surpassing the existing methods across most datasets, showing an average accuracy improvement of 1.08% over DPE. Especially on the fine-grained Flowers102 [Flower] datasaet, our method outperforms the second-best AWT by over 2.3%. We also evaluate our method with SOTA methods on the ImageNet [ImageNet] and its 4 out-of-distribution (OOD) variants, as shown in Table 2. The experimental results demonstrate that our method exhibits competitive generalization performance across all methods. Specifically, our ProtoMM outperforms TDA by 2.58% and AWT by 0.71% on average when using the RN50 backbone. Moreover, with the ViT-B/16 backbone, our method surpasses the second-best method, DOTA, by an average of 1.17% on the OOD dataset. Even compared to the training-based DPE, it achieves an average improvement of 1.03%. These results demonstrate the adaptability of our method to varied distributions. Notably, this generalization is achieved without any test-time optimization or tuning.
@@ 4.3 Ablation Study
Effectiveness of using multimodal prototypes. We examine the contributions of the distributed feature construction module and the multimodal prototype learning module, respectively. The results are summarized in Table 3. Row (1) presents the zero-shot CLIP [CLIP] performance as a baseline. In Row (2), Eq. (5) stands for the distributed feature construction, which expands the textual prototypes by incorporating additional features with their importance weights, achieving 1% gains over CLIP on ImageNet with RN50. Row (3) shows the results of using both modules in our method, outperforming the variant that only uses Eq. (5) by over 1% on Caltech101 with ViT-B/16. These improvements indicate that dynamically incorporating multimodal prototypes leads to better representations and improved test-time adaptation.
@@
Comparison of Inference Time Our method is training-free, unlike methods that require prompt tuning or adapter tuning, such as TPT, DiffTPT, and DPE, which involve backpropagating through an expensive encoder or adapter during optimization. We evaluate the inference time using a single NVIDIA 3090 GPU. The experimental results in Table 4 indicate that our method is faster than the tuning-based methods. Although our method takes slightly more time than TDA, we achieve a notable performance improvement, outperforming TDA by over 2.4%. It is worth noting that the Sinkhorn algorithm [Cuturi13] supports efficient parallel computation, allowing the distance between one distribution and multiple others to be evaluated simultaneously, as demonstrated in prior works [PLOT, AWT]. This parallelism further contributes to the comparable runtime of our method relative to TDA.
@@
CLIP [CLIP] 13min 59.81 0
TPT [TPT] 14h 30min 60.74 +0.93
DiffTPT [DiffTPT] 37h 25min 60.80 +0.99
TDA [TDA] 17min 61.35 +1.54
DOTA [dota] 18min 61.82 +2.01
DPE [DPE] 2h 40min 63.41 +3.60
AWT [AWT] 21min 63.01 +3.20
ProtoMM 21min 63.76 +3.95
@@ 5 Conclusion
6 Acknowledge
This research is supported by the National Natural Science Foundation of China (No.U24B20180) and the advanced computing resources provided by the Supercomputing Center of the USTC. It is also supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR.